### v1

@@ Abstract
We devise and analyse a novel $\boldsymbol{H}(\textbf{curl})$-reconstruction operator for piecewise polynomial fields on shape-regular simplicial meshes. The (non-polynomial) reconstruction is devised over the mesh vertex patches using the partition of unity induced by hat basis functions in combination with local Helmholtz decompositions. Our main focus is on homogeneous tangential boundary conditions. We prove that the difference between the reconstructed $\boldsymbol{H}_{0}(\textbf{curl})$-field and the original, piecewise polynomial field, measured in the broken curl norm and in the $\boldsymbol{L}^{2}$-norm, can be bounded in terms of suitable jump norms of the original field. The bounds are always $h$-optimal, and $p$-suboptimal by $\frac{1}{2}$-order for the broken curl norm and by $\frac{3}{2}$-order for the $\boldsymbol{L}^{2}$-norm. An auxiliary result of independent interest is a novel broken-curl, divergence-preserving Poincaré inequality on vertex patches. Moreover, the $\boldsymbol{L}^{2}$-norm estimate can be improved to $\frac{1}{2}$-order suboptimality under a (reasonable) assumption on the uniform elliptic regularity pickup for a Poisson problem with Neumann conditions over the vertex patches. We also discuss extensions of the $\boldsymbol{H}_{0}(\textbf{curl})$-reconstruction operator to the prescription of mixed boundary conditions, to agglomerated polytopal meshes, and to convex domains. Finally, we showcase an important application of the $\boldsymbol{H}(\textbf{curl})$-reconstruction operator to the $hp$-a posteriori nonconforming error analysis of Maxwell’s equations. We focus on the (symmetric) interior penalty discontinuous Galerkin (dG) approximation of some simplified forms of Maxwell’s equations.
@@ 1 Introduction
We consider the (symmetric) interior penalty discontinuous Galerkin (dG) method for space discretization. This approximation technique is classical for elliptic problems (see, e.g., [3, 18] and the references therein) and has been developed and analyzed in the context of Maxwell’s equations in [36, 29]. Here, we perform the $hp$-a posteriori error analysis hinging on the present ${\boldsymbol{H}}(\text{\bf curl})$}}-reconstruction operator, focusing on homogeneous Dirichlet boundary conditions. This analysis offers, to our knowledge, two ground-breaking features: it is the first time that a full $hp$-analysis is achieved, and it is the first time that the analysis hinges on local (vertex patch) Helmholtz decompositions to estimate the nonconforming error. We emphasize that, while the theoretical derivation of the error estimate relies on the ${\boldsymbol{H}}_{0}(\text{\bf curl})$}}-reconstruction operator, this reconstruction does not need to be computed in practice. Previous a posteriori error estimates bounded the nonconforming error using an averaging operator, as in [30, 31, 12], and were therefore only $h$-optimal. We notice, in passing, that [12] derived an $hp$-a posteriori bound on the conforming error, and addressed the frequency-dependence of the a posteriori estimates in the indefinite regime by using a dualtiy argument à la Schatz.
@@ 2.2 Mesh and polynomial spaces
The set of mesh vertices is denoted by $\mathcal{V}_{h}$ and is decomposed into the subset of interior vertices, $\mathcal{V}_{h}^{\mathrm{int}}$, and the subset of boundary vertices, $\mathcal{V}_{h}^{\mathrm{bd}}$. For all ${\mathbf{a}}\in\mathcal{V}_{h}$, ${\cal{T}}_{{\mathbf{a}}}$ denotes the collection of the mesh cells which share $\mathbf{a}$} and $D_{{\mathbf{a}}}$ the corresponding open subdomain (often called vertex patch). In addition, $\mathcal{F}_{{\mathbf{a}}}$ denotes the collection of the mesh faces which share $\mathbf{a}$}, and $\mathcal{F}_{{\mathbf{a}}}^{\partial}$ denotes the collection of all the mesh faces lying on $\partial D_{{\mathbf{a}}}$ and not containing $\mathbf{a}$}; thus the two sets $\mathcal{F}_{{\mathbf{a}}}$ and $\mathcal{F}_{{\mathbf{a}}}^{\partial}$ are disjoint, and $\mathcal{F}_{{\mathbf{a}}}^{\partial}=\emptyset$ whenever ${\mathbf{a}}\in\mathcal{V}_{h}^{\mathrm{bd}}$. For all ${\mathbf{a}}\in\mathcal{V}_{h}$, we set $h_{\mathbf{a}}:=\operatorname{diam}(D_{{\mathbf{a}}})$, and let $\psi_{{\mathbf{a}}}$ be the hat basis function equal to $1$ at $\mathbf{a}$} and having a support in $D_{{\mathbf{a}}}$. Recall that the hat basis functions satisfy the following partition-of-unity property:
@@ 3 Main results on ${\boldsymbol{H}}(\text{\bf curl})$}}-reconstruction
The ${\boldsymbol{L}}^{2}$-estimate from Lemma 3.4 can be improved by invoking the elliptic regularity pickup of the Poisson problem with homogeneous Neumann boundary conditions on the vertex patches. We need, however, a uniform estimate of the underlying constants over all the vertex patches in the mesh. It is natural to expect that these constants can be bounded in terms of the mesh shape-regularity. As we did not find a precise result in the literature supporting this claim, we formulate it as an assumption. We hope that this will stimulate further research in this direction to prove (or disprove) the assumption.
@@
$\|\boldsymbol{R}^{\mathrm{c}}({\boldsymbol{E}}_{h})-{\boldsymbol{E}}_{h}\|\leq C(\kappa_{\mathcal{T}_{h}})p\bigg{\{}\!\sum_{K\in\mathcal{T}_{h}}\!\Big{(}\frac{h_{K}}{p}\Big{)}^{3}\!\|{[\![{\nabla_{h}{\times}{\boldsymbol{E}}_{h}}]\!]_{\partial K}^{\mathrm{d}}}\|_{\partial K}^{2}\!+\!\!\sum_{K\in\mathcal{T}_{h}}\!h_{K}\|{[\![{{\boldsymbol{E}}_{h}}]\!]_{\partial K}^{\mathrm{c}}}\|_{\partial K}^{2}\!\bigg{\}}^{\frac{1}{2}}\!\!.$ (3.14b)
@@
The bounds (3.14) readily follow from the local estimates established in Lemmas 3.3, 3.4, and 3.6 once we observe that, for all $K\in\mathcal{T}_{h}$, $(\boldsymbol{R}^{\mathrm{c}}({\boldsymbol{E}}_{h})-{\boldsymbol{E}}_{h})|_{K}=\sum_{{\mathbf{a}}\in\mathcal{V}_{K}}{\boldsymbol{\delta}}_{{\mathbf{a}}}|_{K}$ owing to the partition-of-unity property (2.2). Invoking the triangle inequality and the mesh shape-regularity concludes the proof. ∎
@@
where ${\mathcal{F}}_{h}^{\rm N}$ denotes the subset of mesh boundary faces allocated on $\Gamma_{\rm N}$. The local ${\boldsymbol{L}}^{2}$-error estimate relies on the broken Poincaré with mixed boundary conditions discussed in Remark 4.5 below (which requires one more conjecture, unless pure Neumann conditions are enforced). The estimate reads as follows:
@@ 4 Preparatory results
For completeness, we sketch a short proof. We solve the Neumann problem $(\nabla\xi,\nabla w)_{D_{{\mathbf{a}}}}=({\boldsymbol{v}},\nabla w)_{D_{{\mathbf{a}}}}$ in $X^{\mathrm{g}}(D_{{\mathbf{a}}})$, and observe that $\nabla\xi-{\boldsymbol{v}}\in{{{\boldsymbol{H}}_{0}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$. Using [28, Theorem 3], we infer that there exists ${\boldsymbol{\zeta}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ such that $\nabla_{\!0}{\times}{\boldsymbol{\zeta}}=\nabla\xi-{\boldsymbol{v}}.$ Then, we solve the Dirichlet problem $(\nabla\gamma,\nabla w)_{D_{{\mathbf{a}}}}=({\boldsymbol{\zeta}},\nabla w)_{D_{{\mathbf{a}}}}$ in $X^{\mathrm{g}}_{0}(D_{{\mathbf{a}}})$ and set ${\boldsymbol{\psi}}={\boldsymbol{\zeta}}-\nabla\gamma$. The pair $({\boldsymbol{\psi}},\xi)$ satisfies all the required properties. ∎
@@
There exists a constant $C(\kappa_{\mathcal{T}_{h}})$ so that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$ and all ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}}):=\{{\boldsymbol{w}}\in{\boldsymbol{L}}^{2}(D_{{\mathbf{a}}}){\;|\;}\nabla_{h}{\boldsymbol{w}}\in{\boldsymbol{L}}^{2}(D_{{\mathbf{a}}})\}$, there exists ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ satisfying ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$ such that
$\|{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\|_{D_{{\mathbf{a}}}}\!\leq\!{}C(\kappa_{\mathcal{T}_{h}})h_{{\mathbf{a}}}\!\bigg{\{}\!\!\|\nabla_{h}{\times}({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}})\|_{D_{{\mathbf{a}}}}^{2}\!\!+\!\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}+\!\!\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!h_{F}^{-1}\|{\gamma_{K(F),F}^{\mathrm{c}}({\boldsymbol{v}})}\|_{F}^{2}\!\bigg{\}}^{\frac{1}{2}}\!\!,$ (4.2)
where $K(F)$ is the unique mesh cell in ${\cal{T}}_{{\mathbf{a}}}$ of which $F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}$ is a face.
@@
Let ${\mathbf{a}}\in\mathcal{V}_{h}$ and let ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$. Recall that $X^{\mathrm{g}}_{0}(D_{{\mathbf{a}}}):=H^{1}_{0}(D_{{\mathbf{a}}})$ and $X^{\mathrm{g}}(D_{{\mathbf{a}}}):=H^{1}(D_{{\mathbf{a}}})/{\mathbb{R}}$.
(1) Let $\rho_{{\mathbf{a}}}\in X^{\mathrm{g}}_{0}(D_{{\mathbf{a}}})$ solve the following well-posed problem:
$(\nabla\rho_{{\mathbf{a}}},\nabla w_{{\mathbf{a}}})_{D_{{\mathbf{a}}}}=({\boldsymbol{v}},\nabla w_{{\mathbf{a}}})_{D_{{\mathbf{a}}}}\quad\forall w_{{\mathbf{a}}}\in X^{\mathrm{g}}_{0}(D_{{\mathbf{a}}}),$ (4.3)
so that $\nabla\rho_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$. Setting ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}:=\nabla\rho_{{\mathbf{a}}}+\Theta_{{\mathbf{a}}}$ for any $\Theta_{{\mathbf{a}}}\in{\boldsymbol{X}^{\mathrm{c}}_{0}}(D_{{\mathbf{a}}})$, we obtain ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ with ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$.
(2) Since $D_{{\mathbf{a}}}$ is a simply connected, Lipschitz domain, we infer from [2, Lemma 3.5] that there exists ${\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}}):=[X^{\mathrm{g}}(D_{{\mathbf{a}}})]^{d}$ such that $\nabla{\times}{\boldsymbol{\phi}}={\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}$ in $D_{{\mathbf{a}}}$ (notice that $\nabla{\cdot}{\boldsymbol{\phi}}=0$ in $D_{{\mathbf{a}}}$ as well, but this additional property is not needed here). However, the stability constant on $\|{\nabla{\boldsymbol{\phi}}}\|_{D_{{\mathbf{a}}}}$ is not known explicitly in terms of the geometry of $D_{{\mathbf{a}}}$. Therefore, we invoke instead [27, Theorem 35] which gives ${\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}})$ such that $\nabla{\times}{\boldsymbol{\phi}}={\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}$ in $D_{{\mathbf{a}}}$ and a stability constant $C(\kappa_{\mathcal{T}_{h}})$ only depending on the mesh shape-regularity so that the following bound holds:
@@
$\begin{split}\|{{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}}\|_{D_{{\mathbf{a}}}}^{2}&=({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}},\nabla{\times}{\boldsymbol{\phi}})_{D_{{\mathbf{a}}}}\\ &=({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}},\nabla{\times}{\boldsymbol{\phi}})_{D_{{\mathbf{a}}}}-({\boldsymbol{v}},\nabla{\times}{\boldsymbol{\phi}})_{D_{{\mathbf{a}}}}\\ &=-(\nabla_{h}{\times}({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}),\tilde{{\boldsymbol{\phi}}})_{D_{{\mathbf{a}}}}-\sum_{K\in{\cal{T}}_{{\mathbf{a}}}}({\boldsymbol{v}},{\boldsymbol{n}}_{K}\times{\boldsymbol{\phi}})_{\partial K}=:T_{1}+T_{2},\end{split}$
@@
(3a) Bound on $T_{1}$. Using the Cauchy–Schwarz inequality, recalling that $\boldsymbol{\phi}$} has zero mean-value over $D_{{\mathbf{a}}}$, and invoking the Poincaré inequality on the vertex patch (see [38, Theorem 3.2] and also [21, Lem. 5.7]), we infer that
@@
(3b) Bound on $T_{2}$. Re-organizing the summation over the mesh faces in $\mathcal{F}_{{\mathbf{a}}}$ and $\mathcal{F}_{{\mathbf{a}}}^{\partial}$ (recall that these two sets are disjoint) and observing that $\boldsymbol{\phi}$} has continuous traces over all the faces in $\mathcal{F}_{{\mathbf{a}}}$ gives
$T_{2}=\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}}\!([\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F},{\boldsymbol{n}}_{F}\times{\boldsymbol{\phi}})_{F}+\!\sum_{F\in\mathcal{F}_{{\mathbf{a}}}^{\partial}}\!(\gamma_{K(F),F}^{\mathrm{c}}({\boldsymbol{v}}),{\boldsymbol{n}}_{K(F)}\times{\boldsymbol{\phi}})_{F}.$

### v2

@@ Abstract
We devise and analyze a novel ${\boldsymbol{H}}(\text{\bf curl})$}}-reconstruction operator for piecewise polynomial fields on shape-regular simplicial meshes. The (non-polynomial) reconstruction is devised over the mesh vertex patches using the partition of unity induced by hat basis functions in combination with local Helmholtz decompositions. Our main focus is on homogeneous tangential boundary conditions. We prove that the difference between the reconstructed ${\boldsymbol{H}}_{0}(\text{\bf curl})$}}-field and the original, piecewise polynomial field, measured in the broken curl norm and in the ${\boldsymbol{L}}^{2}$-norm, can be bounded in terms of suitable jump norms of the original field. The bounds are always $h$-optimal, and $p$-suboptimal by $\frac{1}{2}$-order for the broken curl norm and by $\frac{3}{2}$-order for the ${\boldsymbol{L}}^{2}$-norm. An auxiliary result of independent interest is a novel broken-curl, divergence-preserving Poincaré inequality on vertex patches. Moreover, the ${\boldsymbol{L}}^{2}$-norm estimate can be improved to $\frac{1}{2}$-order suboptimality under a (reasonable) assumption on the uniform elliptic regularity pickup for a Poisson problem with Neumann conditions over the vertex patches. We also discuss extensions of the ${\boldsymbol{H}}_{0}(\text{\bf curl})$}}-reconstruction operator to the prescription of mixed boundary conditions, to agglomerated polytopal meshes, and to convex domains. Finally, we showcase an important application of the ${\boldsymbol{H}}(\text{\bf curl})$}}-reconstruction operator to the $hp$-a posteriori nonconforming error analysis of Maxwell’s equations. We focus on the (symmetric) interior penalty discontinuous Galerkin (dG) approximation of some simplified forms of Maxwell’s equations.
@@ 1 Introduction
We consider the (symmetric) interior penalty discontinuous Galerkin (dG) method for space discretization. This approximation technique is classical for elliptic problems (see, e.g., [3, 19] and the references therein) and has been developed and analyzed in the context of Maxwell’s equations in [37, 30]. Here, we perform the $hp$-a posteriori error analysis hinging on the present ${\boldsymbol{H}}(\text{\bf curl})$}}-reconstruction operator, focusing on homogeneous Dirichlet boundary conditions. This analysis offers, to our knowledge, two ground-breaking features: it is the first time that a full $hp$-analysis is achieved, and it is the first time that the analysis hinges on local (vertex patch) Helmholtz decompositions to estimate the nonconforming error. We emphasize that, while the theoretical derivation of the error estimate relies on the ${\boldsymbol{H}}_{0}(\text{\bf curl})$}}-reconstruction operator, this reconstruction does not need to be computed in practice. Previous a posteriori error estimates bounded the nonconforming error using an averaging operator, as in [31, 32, 12], and were therefore only $h$-optimal. We notice, in passing, that [12] derived an $hp$-a posteriori bound on the conforming error, and addressed the frequency-dependence of the a posteriori estimates in the indefinite regime by using a dualtiy argument à la Schatz (see [38]).
@@ 2.2 Mesh and polynomial spaces
The set of mesh vertices is denoted by $\mathcal{V}_{h}$ and is decomposed into the subset of interior vertices, $\mathcal{V}_{h}^{\mathrm{int}}$, and the subset of boundary vertices, $\mathcal{V}_{h}^{\mathrm{bd}}$. For all ${\mathbf{a}}\in\mathcal{V}_{h}$, ${\cal{T}}_{{\mathbf{a}}}$ denotes the collection of the mesh cells which share $\mathbf{a}$} and $D_{{\mathbf{a}}}$ the corresponding open subdomain (often called vertex patch). The set of mesh faces belonging to $\overline{D_{{\mathbf{a}}}}$, say $\overline{\mathcal{F}}_{{\boldsymbol{a}}}$}, is partitioned as ${\overline{\mathcal{F}}_{{\boldsymbol{a}}}}=\mathcal{F}_{{\mathbf{a}}}\cup{\mathcal{F}}_{\setminus{\boldsymbol{a}}}$, where $\mathcal{F}_{{\mathbf{a}}}$ is the collection of the mesh faces which share $\mathbf{a}$} and ${\mathcal{F}}_{\setminus{\boldsymbol{a}}}$ the collection of all the mesh faces lying on $\partial D_{{\mathbf{a}}}$ and not containing $\mathbf{a}$}. We notice that the two sets $\mathcal{F}_{{\mathbf{a}}}$ and ${\mathcal{F}}_{\setminus{\boldsymbol{a}}}$ are disjoint and that the set $\mathcal{F}_{{\mathbf{a}}}$ contains mesh boundary faces if $\mathbf{a}$} lies on the boundary. Furthermore, for all ${\mathbf{a}}\in\mathcal{V}_{h}$, we set $h_{\mathbf{a}}:=\operatorname{diam}(D_{{\mathbf{a}}})$, and let $\psi_{{\mathbf{a}}}$ be the hat basis function equal to $1$ at $\mathbf{a}$} and having support in $D_{{\mathbf{a}}}$. Recall that the hat basis functions satisfy the following partition-of-unity property:
@@ 3 Main results on ${\boldsymbol{H}}(\text{\bf curl})$}}-reconstruction
The ${\boldsymbol{L}}^{2}$-estimate from Lemma 3.8 can be improved by invoking the elliptic regularity pickup of the Poisson problem with homogeneous Neumann boundary conditions on the vertex patches. The improvement consists in removing the broken curl contribution from (3.11) and also the global scaling factor $p$ in front of the opening brace. We need, however, a uniform estimate of the underlying constants over all the vertex patches in the mesh. It is natural to expect that these constants can be bounded in terms of the mesh shape-regularity. As we did not find a precise result in the literature supporting this claim, we formulate it as an assumption. We hope that this will stimulate further research in this direction to prove (or disprove) the assumption.
@@
$\|\boldsymbol{R}^{\mathrm{c}}({\boldsymbol{E}}_{h})-{\boldsymbol{E}}_{h}\|\leq C(\kappa_{\mathcal{T}_{h}})p\bigg\{\!\sum_{K\in\mathcal{T}_{h}}\!\Big(\frac{h_{K}}{p}\Big)^{3}\!\|{[\![{\nabla_{h}{\times}{\boldsymbol{E}}_{h}}]\!]_{\partial K}^{\mathrm{d}}}\|_{\partial K}^{2}\!+h_{K}\|{[\![{{\boldsymbol{E}}_{h}}]\!]_{\partial K}^{\mathrm{c}}}\|_{\partial K}^{2}\!\bigg\}^{\frac{1}{2}}\!\!.$ (3.14b)
@@
The bounds (3.14) readily follow from the local estimates established in Lemmas 3.7, 3.8, and 3.10 once we observe that, for all $K\in\mathcal{T}_{h}$, $(\boldsymbol{R}^{\mathrm{c}}({\boldsymbol{E}}_{h})-{\boldsymbol{E}}_{h})|_{K}=\sum_{{\mathbf{a}}\in\mathcal{V}_{K}}{\boldsymbol{\delta}}_{{\mathbf{a}}}|_{K}$ owing to the partition-of-unity property (2.2). Invoking the triangle inequality and the mesh shape-regularity concludes the proof. ∎
@@
where ${\mathcal{F}}_{h}^{\rm N}$ denotes the subset of mesh boundary faces lying on $\Gamma_{\rm N}$. The local ${\boldsymbol{L}}^{2}$-error estimate relies on the broken Poincaré with mixed boundary conditions discussed in Remark 4.21 below (which requires one more conjecture, unless pure Neumann conditions are enforced). The estimate reads as follows:
@@ 4 Preparatory results
For completeness, we sketch a short proof. We solve the Neumann problem $(\nabla\xi,\nabla w)_{D_{{\mathbf{a}}}}=({\boldsymbol{v}},\nabla w)_{D_{{\mathbf{a}}}}$ in $X^{\mathrm{g}}(D_{{\mathbf{a}}})$, and observe that $\nabla\xi-{\boldsymbol{v}}\in{{{\boldsymbol{H}}_{0}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$. Using [29, Theorem 3], we infer that there exists ${\boldsymbol{\zeta}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ such that $\nabla_{\!0}{\times}{\boldsymbol{\zeta}}={{\boldsymbol{v}}-\nabla\xi}.$ Then, we solve the Dirichlet problem $(\nabla\gamma,\nabla w)_{D_{{\mathbf{a}}}}=({\boldsymbol{\zeta}},\nabla w)_{D_{{\mathbf{a}}}}$ in $X^{\mathrm{g}}_{0}(D_{{\mathbf{a}}})$ and set ${\boldsymbol{\psi}}={\boldsymbol{\zeta}}-\nabla\gamma$. The pair $({\boldsymbol{\psi}},\xi)$ satisfies all the required properties. ∎
@@
There exists a constant $C(\kappa_{\mathcal{T}_{h}})$ so that, for all ${\mathbf{a}}\in\mathcal{V}_{h}$, all ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}}):=\{{\boldsymbol{w}}\in{\boldsymbol{L}}^{2}(D_{{\mathbf{a}}}){\;|\;}\nabla_{h}{\boldsymbol{w}}\in{\boldsymbol{L}}^{2}(D_{{\mathbf{a}}})\}$, and all ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ with ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$, the following holds:
$\|{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\|_{D_{{\mathbf{a}}}}\leq{}C(\kappa_{\mathcal{T}_{h}})h_{{\mathbf{a}}}\bigg\{\|\nabla_{h}{\times}({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}})\|_{D_{{\mathbf{a}}}}^{2}+\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!h_{F}^{-1}\|{[\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F}}\|_{F}^{2}\bigg\}^{\frac{1}{2}},$ (4.2)
recalling that the jump operator at $F$ denotes the actual trace whenever $F\in{\mathcal{F}}_{h}^{\mathrm{bd}}$.
@@
Let ${\mathbf{a}}\in\mathcal{V}_{h}$, let ${\boldsymbol{v}}\in{\boldsymbol{H}}^{1}({\cal{T}}_{{\mathbf{a}}})$, and let ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}\in{{{\boldsymbol{H}}_{0}(\text{\bf curl};D_{{\mathbf{a}}})}}$ be such that ${\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}\in{{{\boldsymbol{H}}(\text{\rm div}=0;D_{{\mathbf{a}}})}}$. Recall that $X^{\mathrm{g}}(D_{{\mathbf{a}}}):=H^{1}(D_{{\mathbf{a}}})/{\mathbb{R}}$.
(1) Since $D_{{\mathbf{a}}}$ is a simply connected, Lipschitz domain, we infer from [2, Lemma 3.5] that there exists ${\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}}):=[X^{\mathrm{g}}(D_{{\mathbf{a}}})]^{d}$ such that $\nabla{\times}{\boldsymbol{\phi}}={\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}$ in $D_{{\mathbf{a}}}$ (notice that $\nabla{\cdot}{\boldsymbol{\phi}}=0$ in $D_{{\mathbf{a}}}$ as well, but this additional property is not needed here). However, the stability constant on $\|{\nabla{\boldsymbol{\phi}}}\|_{D_{{\mathbf{a}}}}$ is not known explicitly in terms of the geometry of $D_{{\mathbf{a}}}$. Therefore, we invoke instead [28, Theorem 35]. To do so, we observe that each vertex patch $D_{{\mathbf{a}}}$ can be viewed as a chain of star-shaped domains, as defined in [28, Section 5]. Indeed, following [27, Appendix B], we can enumerate the simplices in the patch as $\{K_{i}\}_{i=1}^{n}$ so that two successive simplices in the enumeration share a face, say $F_{i}=\partial K_{i}\cap\partial K_{i+1}$. We then define the subdomains $D_{{\mathbf{a}}}^{i}:=K_{i}\cup K_{i+1}$ for all $i\in\{1{:}n{-}1\}$. As a consequence of mesh shape-regularity, each $D_{{\mathbf{a}}}^{i}$ is star-shaped with respect to a $d$-dimensional ball centered at the barycenter of the face $F_{i}$, with radius comparable to $h_{F_{i}}$. Moreover, we have the overlap relation $K_{i+1}=D_{{\mathbf{a}}}^{i}\cap D_{{\mathbf{a}}}^{i+1}$. Invoking [28, Theorem 35] gives ${\boldsymbol{\phi}}\in{\boldsymbol{X}}^{\mathrm{g}}(D_{{\mathbf{a}}})$ such that $\nabla{\times}{\boldsymbol{\phi}}={\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}$ in $D_{{\mathbf{a}}}$ and a stability constant $C(\kappa_{\mathcal{T}_{h}})$ only depending on the mesh shape-regularity so that the following bound holds:
@@
$\begin{split}\|{{\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}}\|_{D_{{\mathbf{a}}}}^{2}&=({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}},\nabla{\times}{\boldsymbol{\phi}})_{D_{{\mathbf{a}}}}\\ &=({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}},\nabla{\times}{\boldsymbol{\phi}})_{D_{{\mathbf{a}}}}-({\boldsymbol{v}},\nabla{\times}{\boldsymbol{\phi}})_{D_{{\mathbf{a}}}}\\ &=-(\nabla_{h}{\times}({\boldsymbol{v}}^{\mathrm{c}}_{{\mathbf{a}}}-{\boldsymbol{v}}),{{\boldsymbol{\phi}}})_{D_{{\mathbf{a}}}}-\sum_{K\in{\cal{T}}_{{\mathbf{a}}}}({\boldsymbol{v}},{\boldsymbol{n}}_{K}\times{\boldsymbol{\phi}})_{\partial K}=:T_{1}+T_{2},\end{split}$
@@
(2a) Bound on $T_{1}$. Using the Cauchy–Schwarz inequality, recalling that $\boldsymbol{\phi}$} has zero mean-value over $D_{{\mathbf{a}}}$, and invoking the Poincaré inequality on the vertex patch (see [40, Theorem 3.2] and also [21, Lem. 5.7]), we infer that
@@
(2b) Bound on $T_{2}$. Re-organizing the summation over the mesh faces in $\overline{\mathcal{F}}_{{\boldsymbol{a}}}$} as those in ${\overline{\mathcal{F}}_{{\boldsymbol{a}}}}\cap{\mathcal{F}}_{h}^{\mathrm{int}}$ where $\boldsymbol{\phi}$} has continuous trace and those in ${\overline{\mathcal{F}}_{{\boldsymbol{a}}}}\cap{\mathcal{F}}_{h}^{\mathrm{bd}}$ where the jump operator denotes the actual trace, we obtain
$T_{2}=\!\sum_{F\in{{\overline{\mathcal{F}}_{{\boldsymbol{a}}}}}}\!([\![{{\boldsymbol{v}}}]\!]^{\mathrm{c}}_{F},{\boldsymbol{n}}_{F}\times{\boldsymbol{\phi}})_{F}.$